# HuMam: Humanoid Motion Control via End-to-End Deep Reinforcement Learning with Mamba

### v1

@@ IV-D2 Effect on Learning Efficiency and Stability
IV-D3 Effect on Reward-Based Performance
Table IV provides a detailed breakdown of the six reward components across tasks. Overall, HuMam achieves higher or comparable scores in nearly all dimensions, leading to an average per-step reward of $0.737$, surpassing the Baseline ($0.728$).
@@ IV-D3 Effect on Reward-Based Performance
The footstep trajectories of different walking modes are shown in Figure 4,  5,  6, and  7. Although Baseline shows marginally higher step-placement accuracy in curved and forward walking, HuMam achieves better balance across all reward terms. This indicates that the Mamba encoder does not over-optimize a single component (e.g., foot placement) at the expense of global stability, but instead integrates multimodal inputs into a more holistic control strategy.
@@
TABLE IV: Comprehensive Reward Components Comparison
@@
IV-D4 Effect on Joint Torque Profiles
Table V and Figure  8 present the torque statistics for the forward walking task. On average, HuMam reduces joint torques by 9.6% in magnitude and 9.1% in peak load, indicating smoother and less strenuous actuation compared with the Baseline.
@@ IV-D4 Effect on Joint Torque Profiles
TABLE V: Joint Torque Comparison of Forward Walking Task
@@
IV-D5 Effect on Energy Efficiency
Table VI reports the energy efficiency, average power, and normalized power per unit mass across tasks. Overall, HuMam demonstrates lower power consumption and improved economy of movement compared with the Baseline.
@@ IV-D5 Effect on Energy Efficiency
These results indicate that the Mamba encoder helps produce smoother joint actuation and avoids unnecessary corrective torques, directly lowering the energy required per unit distance. Combined with the improvements in reward composition (Section IV-D) and reduced joint torque magnitudes (Table V), the evidence consistently supports that HuMam enables more energy-efficient, stable, and sustainable locomotion than conventional feedforward baselines.
TABLE VI: Energy Efficiency and Power Consumption
Task Energy Efficiency Average Power Power Efficiency
(J/m) (W) (W/kg)
@@
Standing - - 35.08 57.48 0.56 0.92
Lateral 963 1,395 95.45 121.10 1.53 1.94
Curved 1,303 1,055 95.17 131.93 1.53 2.11

### v2

@@ 4.4.2 Effect on Learning Efficiency and Stability
4.4.3 Effect on Reward-Based Performance
Table 4 provides a detailed breakdown of the six reward components across tasks. Overall, HuMam achieves higher or comparable scores in nearly all dimensions, leading to an average per-step reward of $0.737$, surpassing the Baseline ($0.728$).
@@ 4.4.3 Effect on Reward-Based Performance
Although Baseline shows marginally higher step-placement accuracy in curved and forward walking, HuMam achieves better balance across all reward terms. This indicates that the Mamba encoder does not over-optimize a single component (e.g., foot placement) at the expense of global stability, but instead integrates multimodal inputs into a more holistic control strategy.
@@
Table 4: Comprehensive Reward Components Comparison
@@
4.4.4 Effect on Joint Torque Profiles
Table 5 and Fig.  8 present the torque statistics for the forward walking task. On average, HuMam reduces joint torques by 9.6% in magnitude and 9.1% in peak load, indicating smoother and less strenuous actuation compared with the Baseline.
@@ 4.4.4 Effect on Joint Torque Profiles
Table 5: Joint Torque Comparison of Forward Walking Task
@@
4.4.5 Effect on Energy Efficiency
Table 6 reports the energy efficiency, average power, and normalized power per unit mass across tasks. Overall, HuMam demonstrates lower power consumption and improved economy of movement compared with the Baseline.
@@ 4.4.5 Effect on Energy Efficiency
These results indicate that the Mamba encoder helps produce smoother joint actuation and avoids unnecessary corrective torques, directly lowering the energy required per unit distance. Combined with the improvements in reward composition (Section 4.4) and reduced joint torque magnitudes (Table 5), the evidence consistently supports that HuMam enables more energy-efficient, stable, and sustainable locomotion than conventional feedforward baselines.
Table 6: Energy Efficiency and Power Consumption
Task Energy Efficiency (J/m) Average Power (W) Power Efficiency (W/kg)
@@
Standing 35.08 57.48 0.56 0.92
Lateral 963 1395 95.45 121.10 1.53 1.94
Curved 1303 1055 95.17 131.93 1.53 2.11
@@ 5 Conclusion
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.